You are using a quadrat to estimate the number of squirrel nests in your neighborhood park. The park is 10 square miles. You divide park into 100 equally sized sections. You count the number of squirrel nests in 10 of the 100 sections. In those 10 areas you count a total of 100 nests.

Based on this information, you estimate the total number of squirrel nests in the park as:
Answer #1

total number of section in 10 square miles = 100

number of total nest in 10 sections =100

number of nest in one section = 100/10 =10

so the total number of nest in park that is 10 section = 10 * 100 = 1000



answered by: ANURANJAN SARSAM